OpenStudy (anonymous):

A professor must randomly select 4 students to participate in a mock debate. There are 15 students in his class. In how many different ways can these students be selected, if the order of selection does not matter? A. 445 B. 989 C. 32,760 D. 1,365

OpenStudy (anonymous):

you are being asked for "15 choose 4" sometimes written as\(_{15}C_4\) or \(^{15}C_4\) but adults write \(\binom{15}{4}\) computed as \[\binom{15}{4}=\frac{15\times 14\times 13\times 12}{4\times 3\times 2}\] cancel first multiply last

OpenStudy (anonymous):

or just use a calculator http://www.wolframalpha.com/input/?i=15+choose+4

OpenStudy (anonymous):

@satellite73 Cool, thank you!

OpenStudy (anonymous):

yw